OpenStudy (bluebeta):

HELP PLEASE Identify the center and radius for the following equations: 1. x^2+y^2-12x+6y+45=0 2. x^2+y^2-6x+14y+9=0 Graph the following circle: 1. (x-1)^2+(y+6)^2=0

OpenStudy (bluebeta):

@dayakar

OpenStudy (bluebeta):

@misty1212

OpenStudy (anonymous):

1. 75 - 152 2. 69 - 173 1. /--\ | | \__/

OpenStudy (trojanpoem):

You have the equation in general form : x^2 + y^2 + ax + by + c = 0 Complete the square to get the standard form of circle equation: (x-a)^2 + (y-b)^2 = r^2 **************************************************************